A capacitor is connected to a 17 kHz oscillator that produces an rms voltage of 9.0 V . The peak current is 65 mA .

What is the value of the capacitance C?
Answer #1

Xc = 1/wC

V = IXc

Irms = Ipeak/sqrt(2)

9 = 65*10^-3 / sqrt(2) *Xc

Xc = 195.814 = 1/wC

C = 1/(2*pi*17*10^3 *195.814)

C = 47.8 nF